# The environment of TeV halo progenitors

### v1

@@ 4.3.2 Pulsars likely travelling in the ISM
J0742-2822 (also named B0740-28): This pulsar sports a clear optical bow-shock (Jones et al. 2002) and as such could be considered as being in the predominantly neutral ISM, away from its parent environment. With $\tau_{\mathrm{c}}=157\leavevmode\nobreak\ \mathrm{kyr}$, PSR J0742-2822 is in the age range where the transition between mostly in to mostly out of the parent environment occurs. Accordingly, our model predicts a relatively broad probability distribution from 0 to 90% (95% confidence interval). The observed stand-off distance favours small inclination of the kick velocity vector with respect to the plane of the sky (Jones et al. 2002), which would in turn favour high probabilities of residing in the parent environment. On the other hand, the morphology of the optical nebula suggests that it recently left a low-density medium ($n_{\mathrm{H}}\lesssim 3\times 10^{-3}\,\mathrm{H}\,\mathrm{cm}^{-3}$) to enter a denser medium ($n_{\mathrm{H}}\simeq 3\times 10^{-1}\,\mathrm{H}\,\mathrm{cm}^{-3}$), which can be interpreted as a transition from the interior of a bubble to the bounding shell of ambient ISM.
PSR J1741–2054: the pulsar displays a clear bow-shock nebula in optical lines (Mignani et al. 2016; Romani et al. 2010) and UV (Abramkin et al. 2025). This strongly suggests the pulsar propagates in the mostly neutral ISM, which is consistent with our model prediction that this pulsar has a median probability of about 95% of being out of its parent environment. Both Camilo et al. (2009) and Abramkin et al. (2025) noticed that the properties of J1741-2054 are very close to that of the ’Three Musketeers’ (Geminga, B0656+14, and B1055-52). Since TeV halos were discovered around the first two, and that B1055-52 is in the blind spot of HAWC, J1741-2054 appears as a promising target to search for a TeV halo and make a comparison with the most established instances of the phenomenon. If a halo were to be discovered around J1741-2054, our model would suggest that pair confinement is linked to some properties of the ISM rather than being due to the parent environment.
@@ 4.3.3 The case of Geminga
With a characteristic age of $342\leavevmode\nobreak\ \mathrm{kyr}$ and a transverse speed of $150\leavevmode\nobreak\ \mathrm{km/s}$, PSR J0633+1746 has travelled at least $\sim 53\leavevmode\nobreak\ \mathrm{pc}$ since its birth. Because of this large distance, it is commonly assumed that Geminga is in the ISM, and our work also tends to statistically place it there, in terms of median probability. This result seems to contrast with a number of arguments put forward previously and according to which our model overall tends to suggest that the parent environment would be statistically favored as the location where halos develop. Out of seven TeV halos and halo candidates in our sample, four have high probabilities $\gtrsim 80\%$ of residing in their parent environment (J0631+1036, B0656+14, J0633+0632, and J0248+6021) while two have about probability $\sim 50\%$ of being in or out (J0622+3749 and B0540+23). But for Geminga, the prototypical pulsar halo, our model would indicate a different picture. Yet, the characteristic age of Geminga corresponds to the time period when most pulsars escape their parent environment, and our model correspondingly predicts a broad probability distribution from 0 to 80% (95% confidence interval). This mitigates the conclusions based on median probability alone. In addition, three considerations can reconcile Geminga with the rest of halos and candidates.
@@
First, as summarized in Amato & Recchia (2024), there are many hints that Geminga is still in a hot environment. The non-detection of Hα lines in the near vicinity of the pulsar suggests that the medium Geminga traverses is hot and highly ionized. The ambient medium density inferred from Geminga’s X-ray jets is quite low, $n_{\mathrm{H}}\lesssim 10^{-2}$ H cm-3, which is typical of a bubble interior. This seems consistent with the fact that Geminga is located close to the Gemini Hα ring, a bubble that hosts an OB association. Using the formalism from Weaver et al. (1977) and comparing with Suzaku observations of the Hα ring, Knies et al. (2018) shows that the Gemini ring can be seen as a bubble blown by at least two stars of an OB association found inside. Fang et al. (2019) noticed that Geminga and the brightest part of its halo are found in projection inside the Gemini Hα ring.
@@
Second, the morphology of X-ray jets around Geminga suggests that the kick velocity of the pulsar is nearly perpendicular to the line of sight (hence in the plane of the sky), especially when jets are interpreted as bent polar outflows (Posselt et al. 2017). In Fig. 6, this implies that the actual position of Geminga is closer to the star symbol, which implies a probability of residing in its parent environment above 80%.
@@
Third and last, the discussion so far is based on the characteristic age of Geminga. As mentioned earlier, the characteristic age may well overestimate the true age by a factor of a few (a caveat valid for all objects in our sample). Since Geminga is close to the age range where most pulsars transitions from inside to outside of their parent environment, any decrease in age has significant consequences. This is illustrated in Fig. 7, where we display box plots similar to Fig. 6 for a variety of age assumptions for Geminga between $150-342\leavevmode\nobreak\ \mathrm{kyr}$. We see that if Geminga’s true age is roughly half of its characteristic age, its median probability of being inside its parent environment reaches $\gtrsim 70\%$.
@@
Figure 7: Probability that Geminga resides in its parent environment for five different assumptions on its age. For each age, a box plot like those in Fig. 6 is displayed.
@@ 5 Conclusions
In this work, we aimed at understanding the type of medium crossed by pulsars over their lifetime in order to constrain the physical mechanism leading to strong confinement of electron-positron pairs and ultimately to the formation of gamma-ray halos observed at $\gtrsim$ TeV photon energies around some pulsars.
We developed a statistical model for the environment probed by pulsars over the first few Myr of their life, under two different scenarios for the circumstellar conditions: pulsars either result from isolated massive star progenitors surrounded by wind-blown bubbles (WBBs) at the time of supernova, or they result from massive stars belonging to clusters that eventually explode inside superbubbles (SBs). We also produced a mixed scenario from a weighted combination of these two cases to describe the population of our Galaxy.
Using a Monte-Carlo method to sample over progenitor properties and pulsar kick velocities, we computed the statistical distribution of the times at which pulsars escape their parent SNR first, and later their parent WBB or SB interior, to eventually enter into the undisturbed ISM. We applied this model to a fully synthetic population of pulsars and to a sample of known pulsars that either have the right properties to produce detectable halos, or were claimed to be halo candidates on the basis of gamma-ray observations. Our main conclusions are:
@@
    Pulsars escape into the ISM much later than commonly assumed. Only half of a Galactic population of synthetic pulsars have left their parent environments at $200\leavevmode\nobreak\ \mathrm{kyr}$, compared to the typical escape times of $40-60\leavevmode\nobreak\ \mathrm{kyr}$ frequently used in the literature.
@@
    In our model, middle-aged pulsars with ages $100-300$ kyr, those presumably involved in TeV halos, have comparable chances of being inside their or outside their parent environment. While this does not favour any explanation for the origin of halos in terms of medium they develop into, this has implications on the occurrence rate of TeV halos: if the parent environment, stellar-wind bubble or superbubble, provides the turbulence for efficient trapping of ultrarelativistic electron-positron pairs, our model predicts a strongly decreasing occurrence rate of TeV halos as pulsars get older than 100 kyr. At the characteristic age of Geminga, only a third of pulsars would then be able to form halos.
@@
    From a sample of seven known pulsars with a confirmed or suspected halo, the majority have a significant probability of still being inside their parent environment. Geminga appears as a notable exception, with a 90% median probability of being in the ISM at face value. Yet, including additional information on the kick velocity orientation and considering the uncertainty of the pulsar age increase the probability that Geminga resides in its parent environment up to above $70-80$%, which agrees with the hot and ionized properties inferred for the medium around Geminga.
We provide the probability of residing inside the parent environment for a list of 53 known pulsars. Searching for extended gamma-ray emission around them with HAWC or LHAASO, and in the future with the upcoming SWGO experiment, could help us connect the occurrence of the pulsar halo phenomenon with the environment, at least from a statistical point of view. Together with complementary studies of the ISM or CSM (for instance searching for hot ionized gas in X-rays, or for neutral gas in optical lines), this should contribute to improving our knowledge of the mechanisms underlying the formation of TeV halos.
Beyond these considerations, the prediction from our model that pulsars escape their parent environment much later than commonly assumed will have some implications on their contribution to the flux of cosmic-ray positrons and electrons directly measured locally. All other things being equal, a delayed escape probably leads to a smaller flux at Earth, because a smaller fraction of the pairs produced by pulsars will be effectively injected in the ISM. Accounting for the AMS-02 positron spectrum thus implies either a higher pair production efficiency, or the contribution from a complementary source class, or different cosmic-ray transport assumptions near and/or far from the sources.

### v2

@@ 4.3.2 Pulsars likely travelling in the ISM
J0742-2822/B0740-28
This pulsar sports a clear optical bow shock (Jones et al. 2002) and, as such, could be considered to be in the predominantly neutral ISM, away from its parent environment. With $\tau_{\mathrm{c}}=157\penalty 10000\ \mathrm{kyr}$, PSR J0742-2822 is in the age range where the transition between mostly in to mostly out of the parent environment just starts to occurs. Accordingly, our model predicts a probability distribution from 85 to 90% (95% confidence interval). The observed stand-off distance favours small inclination of the kick velocity vector with respect to the plane of the sky (Jones et al. 2002), which would in turn favour high probabilities of residing in the parent environment. On the other hand, the morphology of the optical nebula suggests that it recently left a low-density medium ($n_{\mathrm{H}}\lesssim 3\times 10^{-3}\,\mathrm{H}\,\mathrm{cm}^{-3}$) to enter a denser medium ($n_{\mathrm{H}}\simeq 3\times 10^{-1}\,\mathrm{H}\,\mathrm{cm}^{-3}$), which can be interpreted as a transition from the interior of a bubble to the bounding shell of ambient ISM.
PSR J1741–2054
The pulsar displays a clear bow-shock nebula in optical lines (Mignani et al. 2016; Romani et al. 2010) and UV (Abramkin et al. 2025). This strongly suggests the pulsar propagates in the mostly neutral ISM, which is weakly consistent with our model prediction of an approximately median 60% probability that it resides outside its parent environment. Both Camilo et al. (2009) and Abramkin et al. (2025) noticed that the properties of J1741-2054 are very close to that of the ’Three Musketeers’ (Geminga, B0656+14, and B1055-52). Since TeV haloes were discovered around the first two, and that B1055-52 is in the blind spot of HAWC, J1741-2054 appears as a promising target to search for a TeV halo and make a comparison with the most established instances of the phenomenon. If a halo were discovered around J1741-2054, our model would suggest pair confinement is linked to some properties of the ISM rather than be due to the parent environment.
@@ 4.3.3 The case of Geminga
With a characteristic age of $342\penalty 10000\ \mathrm{kyr}$ and a transverse speed of $150\penalty 10000\ \mathrm{km/s}$, PSR J0633+1746 has travelled at least $\sim 53\penalty 10000\ \mathrm{pc}$ since its birth. Because of this large distance, it is commonly assumed that Geminga is in the ISM.
However, our model statistically places the TeV haloes and halo candidates in our sample inside their parent environment, including Geminga. Since Geminga is usually taken as the prototypical pulsar halo, we provide three additional elements that further support the idea that it is not probing the ISM.
@@
First, as summarised in Amato and Recchia (2024), there are many hints that Geminga is still in a hot environment. The non-detection of Hα lines in the near vicinity of the pulsar suggests that the medium Geminga traverses is hot and highly ionised. The ambient medium density inferred from Geminga’s X-ray jets is quite low, $n_{\mathrm{H}}\lesssim 10^{-2}$ H cm-3, which is typical of a bubble interior. This seems consistent with the fact that Geminga is located close to the Gemini Hα ring, a bubble that hosts an OB association. Using the formalism from Weaver et al. (1977) and comparing with Suzaku observations of the Hα ring, Knies et al. (2018) shows that the Gemini ring can be seen as a bubble blown by at least two stars of an OB association found inside. Fang et al. (2019) noticed that Geminga and the brightest part of its halo are found in projection inside the Gemini Hα ring.
@@
Second, the morphology of X-ray jets around Geminga suggests that the kick velocity of the pulsar is nearly perpendicular to the line of sight (hence in the plane of the sky), especially when jets are interpreted as bent polar outflows (Posselt et al. 2017). In Fig. 6, this implies that the actual position of Geminga is closer to the star symbol, which implies a probability of residing in its parent environment above 90%.
@@
Third and last, the discussion so far is based on the characteristic age of Geminga. As mentioned earlier, the characteristic age may well overestimate the true age by a factor of a few (a caveat valid for all objects in our sample). Since Geminga is exactly at the age range where most pulsars transitions from inside to outside of their parent environment, any decrease in age has significant consequences. This is illustrated in Fig. 7, where we display box plots similar to Fig. 6 for a variety of age assumptions for Geminga between $150-342\penalty 10000\ \mathrm{kyr}$. We see that if Geminga’s true age is roughly half of its characteristic age, its median probability of being inside its parent environment reaches $\gtrsim 90\%$.
@@
Figure 7: Probability that Geminga resides in its parent environment for five different assumptions on its age. For each age, a box plot such as those in Fig. 6 is displayed.
@@ 5 Conclusions
In this work, we aimed at understanding the type of medium crossed by pulsars over their lifetime in order to constrain the physical mechanism leading to strong confinement of electron-positron pairs and ultimately to the formation of gamma-ray haloes observed at greater than or approximately equal to TeV photon energies around some pulsars. We developed a statistical model for the environment probed by pulsars over the first few megayears of their life, under two different scenarios for the circumstellar conditions: pulsars resulting from either isolated massive star progenitors surrounded by WBBs at the time of a supernova or massive stars belonging to clusters that eventually explode inside SBs. We also produced a mixed scenario from a weighted combination of these two cases to describe the population of our Galaxy.
Using a Monte-Carlo method to sample over progenitor properties and pulsar kick velocities, we computed the statistical distribution of the times at which pulsars escape their parent SNR first, and later their parent WBB or SB interior, to eventually enter into the undisturbed ISM. We applied this model to a fully synthetic population of pulsars and to a sample of known pulsars that either have the right properties to produce detectable haloes or were claimed to be halo candidates on the basis of gamma-ray observations. Our main conclusions are the following:
@@
    Pulsars escape into the ISM much later than commonly assumed. Only half of a Galactic population of synthetic pulsars have left their parent environments at $300\penalty 10000\ \mathrm{kyr}$, compared to the typical escape times of $40-60\penalty 10000\ \mathrm{kyr}$ frequently used in the literature.
@@
    In our model, middle-aged pulsars with ages 100300100-300 kyr, those presumably involved in TeV haloes, have higher chances of being inside their parent environment. While this does not favour any explanation for the origin of haloes in terms of the medium they develop into, this has implications on the occurrence rate of TeV haloes: if the parent environment, stellar-wind bubble or SB, provides the turbulence for efficient trapping of ultrarelativistic electron-positron pairs, our model predicts a strongly decreasing occurrence rate of TeV haloes as pulsars get older than 300 kyr. At the characteristic age of Geminga, a half of pulsars would then be able to form haloes.
@@
    From a sample of seven known pulsars with a confirmed or suspected halo, all of them have a significant probability of being inside their parent environment. Geminga has the highest median probability in the sample (only $35\%$) of being in the ISM at face value. Yet, including additional information on the kick velocity orientation and considering the uncertainty of the pulsar age increase the probability that Geminga resides in its parent environment to above 9090%, consistent with the hot and ionised properties inferred for the medium around Geminga.
We provide the probability of residing inside the parent environment for a list of 53 known pulsars. Searching for extended gamma-ray emission around them with HAWC, LHAASO, and, in the future, the upcoming SWGO experiment could help us connect the occurrence of the pulsar halo phenomenon with the environment, at least from a statistical point of view. Together with complementary studies of the ISM or circumstellar medium (for instance searching for hot ionised gas in X-rays, or for neutral gas in optical lines), this should contribute to our knowledge of the mechanisms underlying the formation of TeV haloes.
Beyond these considerations, our model’s prediction that pulsars escape their parent environment much later than commonly assumed will have implications for their contribution to the flux of cosmic-ray positrons and electrons directly and locally measured. All else being equal, a delayed escape likely reduces the flux at Earth, as a smaller fraction of the pairs produced by pulsars will be effectively injected in the ISM. Accounting for the AMS-02 positron spectrum thus implies either higher pair production efficiency, contributions from a complementary source class, or different cosmic-ray transport assumptions near and/or far from the sources.
@@ Acknowledgements.
Appendix A Properties of the individual pulsars studied
This table compiles the properties of all the individual pulsars that fit the conditions given in Sec. 4.1. The probability in the second column represents the probability of pulsars without proper motion information being found inside the parent environment, and the median probability (with the $1\sigma$ interval) for pulsars with proper motion information, in both cases corresponding to the combined WBB+SB scenario for a Galactic population. In the last column, we provide the identification as TeV halo or halo candidate, and as PWN for a normal and fast-moving pulsar (PWN and SPWN, respectively). The data shown come from the ATNF catalogue (Manchester et al. 2005) unless a different reference is provided: (1) Khabibullin et al. (2024b); (2) Van Etten et al. (2012); (3) Danilenko et al. (2019); (4) Johnston and Lower (2021); (5) Pletsch et al. (2012).
Table 2: Properties of the known pulsars selected for comparison to our model.
Pulsar name Probability [%] $1\sigma$ interval [%] $\tau_{\mathrm{c}}$ [kyr] $v_{\mathrm{kick}}$ [km/s] $\dot{E}$ [erg/s] $d$ [kpc] Association
B2334+61 99.87 [99.80; 99.90] 40.6 50 $6.30\times 10^{34}$ 0.7 -
J0538+2817 99.86 [96.70; 99.90] 37.6(1) 357 $4.90\times 10^{34}$ 1.3 PWN
J1809-2332 99.85 [94.83; 99.90] 67.6(2) 112 $4.30\times 10^{35}$ 0.88 PWN
J1044-5737 97.79 - 40.3 - $8.00\times 10^{35}$ 1.895 -
J0940-5428 97.48 - 42.2 - $1.90\times 10^{36}$ 0.377 -
J0631+1036 97.24 - 44.4 - $1.70\times 10^{35}$ 2.104 Candidate Halo
J0554+3107 96.39 - 51.7 - $5.60\times 10^{34}$ 2 -
J1429-5911 95.43 - 60.2 - $7.70\times 10^{35}$ 1.955 -
J1459-6053 94.98 - 64.7 - $9.10\times 10^{35}$ 1.84 -
J1954+2836 94.63 - 69.4 - $1.00\times 10^{36}$ 1.96 -
B1916+14 93.51 - 88.1 - $5.10\times 10^{33}$ 1.304 -
J1028-5819 93.39 - 90 - $8.30\times 10^{35}$ 1.423 -
J1732-$31$ 92.41 - 111 - $1.50\times 10^{35}$ 0.641 -
J1740+1000 92.27 - 114 - $2.30\times 10^{35}$ 1.227 PWN
J0633+0632 92.09 [91.31; 97] 59.2(3) 338 $1.20\times 10^{35}$ 1.355(3) Candidate Halo
B0656+14 91.16 [90; 97.10] 111 60 $3.80\times 10^{34}$ 0.286 TeV Halo
J0248+6021 90.00 [90; 90.73] 62.4 644 $2.10\times 10^{35}$ 2 TeV Halo
B1727-47 90.00 [89.91; 90.03] 80.4 501 $1.10\times 10^{34}$ 0.7 -
J2032+4127 89.62 [84.76; 91.29] 201 19 $1.50\times 10^{35}$ 1.33 -
B0740-28 89.54 [87.29; 90] 157 278 $1.40\times 10^{35}$ 2 SPWN
B1822-09 89.19 [78.25; 90.28] 233 22 $4.50\times 10^{33}$ 0.3 -
B0906-49 88.72 [87.09; 89.39] 112 670 $4.90\times 10^{35}$ 3(4) SPWN
J0954-5430 85.26 - 171 - $1.60\times 10^{34}$ 0.433 -
J1301-6310 82.36 - 186 - $7.60\times 10^{33}$ 1.458 -
B0114+58 82.21 [46.60; 89.76] 275 154 $2.20\times 10^{35}$ 1.768 -
B0941-56 80.76 [28.39; 89.79] 323 33 $3.00\times 10^{33}$ 0.406 -
J0622+3749 77.95 - 208(5) - $2.70\times 10^{34}$ 1.6(5) TeV Halo
J0905-5127 74.97 - 221 - $2.40\times 10^{34}$ 1.33 -
B0540+23 68.04 - 253 - $4.10\times 10^{34}$ 1.565 Candidate Halo
J0633+1746 64.50 [13.19; 87.16] 342 152 $3.20\times 10^{34}$ 0.19 TeV Halo, SPWN
J1549-$48$ 54.26 - 324 - $2.30\times 10^{34}$ 1.308 -
B0922-52 52.24 - 335 - $3.40\times 10^{33}$ 0.513 -
J1846+0919 48.28 - 360 - $3.40\times 10^{34}$ 1.53 -
J0426+4933 46.82 - 371 - $2.00\times 10^{33}$ 1.795 -
J1741-2054 40.16 [6.28; 82.77] 386 155 $9.50\times 10^{33}$ 0.3 SPWN
B0959-54 38.51 - 441 - $6.90\times 10^{32}$ 0.3 -
B0940-55 36.52 - 464 - $3.10\times 10^{33}$ 0.3 -
J1849-0317 34.76 - 481 - $2.90\times 10^{33}$ 1.212 -
J2030+4415 29.69 - 555 - $2.20\times 10^{34}$ 0.72 -
J2017-2737 28.55 - 583 - $2.10\times 10^{34}$ 1.549 -
J1957+5033 21.55 - 838 - $5.30\times 10^{33}$ 1.365 -
B0727-18 21.22 [2.97; 74.07] 426 177 $5.60\times 10^{33}$ 2 -
J0835-3707 20.94 - 877 - $2.40\times 10^{33}$ 0.555 -
J1152-6012 20.69 - 893 - $4.90\times 10^{33}$ 1.367 -
J0212+5222 20.58 - 904 - $4.90\times 10^{33}$ 1.558 -
J1848+0647 20.43 - 916 - $2.70\times 10^{33}$ 1.128 -
J1530-5327 19.95 - 944 - $8.50\times 10^{33}$ 1.123 -
B1055-52 17.66 [1.10; 86.20] 535 21 $3.00\times 10^{34}$ 0.093 -
B1742-30 11.36 [0.71; 82.23] 546 47 $8.50\times 10^{33}$ 0.2 -
B0355+54 9.20 [0.55; 77.19] 564 59 $4.50\times 10^{34}$ 1 SPWN
B0919+06 0.63 [<0.001; 2.35] 497 461 $6.80\times 10^{33}$ 1.1 -
J0357+3205 <0.001 [<0.001; <0.001] 540 649 $5.90\times 10^{33}$ 0.835 SWPN
B1133-55 <0.001 [<0.001; 0.01] 702 318 $6.70\times 10^{33}$ 1.518 -
Table 2: continued.